McConnell Corp. has a book net worth of $13,211. Long-term debt is $2,573. Net working capital, other than cash, is $3,462. Fixed assets are $2,705 and current liabilities are $1,118.

How much cash does the firm have?
Answer #1

total liabilities = Book net worth + Long term debt + current liabilities

= 13211 + 2573 + 1118

= 16902

total assets excluding cash = fixed assets + net working capital + current liabilities

= 2705 + 3462 + 1118

= 7285

cash = total assets excluding cash - total liabilities

= 16902 - 7285

= 9617